# Nuclear density dependence of polarization transfer in quasi-elastic ${\rm A}(\vec{e},e^{\prime}\vec{p})$ reactions

### v2

@@ Abstract
The ratio of the transverse and longitudinal component of polarization transfer to protons in quasi-elastic $(\vec{e},e^{\prime}\vec{p}\,)$ reaction, $P^{\prime}_{x}/P^{\prime}_{z}$, is sensitive to the proton’s electromagnetic form factor ratio, $G_{E}/G_{M}$. To explore density-dependent in-medium modifications, a comparison of polarization transfer ratios involving protons from distinct nuclear shells, each with different local nuclear densities, has been proposed. In this study, we present such comparisons between four shells, $1s_{1/2}$, $1p_{3/2}$ in ${}^{12}\mathrm{C}$ and $1d_{3/2}$, $2s_{1/2}$ in ${}^{40}\mathrm{Ca}$. In an effort to account for other many-body effects that may differ between shells, we use state-of-the-art relativistic distorted-wave impulse-approximation (RDWIA) calculation and present the double ratios, $(P^{\prime}_{x}/P^{\prime}_{z})_{\rm Data}/(P^{\prime}_{x}/P^{\prime}_{z})_{\rm RDWIA}$ as well as the super ratios, $\left[(P^{\prime}_{x}/P^{\prime}_{z})_{\rm A}/(P^{\prime}_{x}/P^{\prime}_{z})_{\rm B}\right]_{\rm Data}/\left[(P^{\prime}_{x}/P^{\prime}_{z})_{\rm A}/(P^{\prime}_{x}/P^{\prime}_{z})_{\rm B}\right]_{\rm RDWIA}$, for chosen shells A and B, as a function of effective local nuclear densities. We find that double ratios for individual shells show a dependence on the probed effective nuclear densities. Studying the super ratios, we observed a systematic variation between pairs of higher- and lower-density shells.
@@ I Introduction
Polarization transfer to a proton bound in a nucleus has been suggested as a tool to observe in-medium modifications in the bound proton structure [1]. It is a part of a wider effort to understand role of quarks and gluons in nuclei [2]. Some calculations introduce in-medium modifications and suggest nuclear-density-dependent changes of the bound nucleon electromagnetic (EM) form factors (FFs) [3, 4, 5]. We report on the first systematic search for nuclear density-dependent effects in the quasi-elastic $A(\vec{e},e^{\prime}\vec{p})$ reaction, which is sensitive to EM FFs.
@@
Density-dependent modifications are expected to be at the level of a few percent [5], which requires high statistical accuracy. ${}^{12}\mathrm{C}$ was suggested as a good nucleus for such studies since the effective local nuclear density experienced by the protons bound in the $1s_{1/2}$ shell is about twice the density for those in the $1p_{3/2}$ shell. Comparison of polarization ratios for protons from $1s_{1/2}$ and $1p_{1/2}$ shells in ${}^{12}\mathrm{C}$ has shown that the $p$ to $s$ double ratio was $1.15\pm 0.03$ [6]. Relativistic distorted-wave impulse-approximation (RDWIA) calculations only partially accounted for this deviation from unity, while relativistic plane-wave impulse-approximation (RPWIA) did not predict this difference. However, once protons were compared at the same virtuality, the results became consistent with unity, $1.05\pm 0.05$, and so did RPWIA and RDWIA calculations [6]. We note that as seen in Fig. 2 of Ref. [6], in the virtuality overlap region, we are effectively comparing the low-$p_{\rm miss}$ region of $1s_{1/2}$ to high-$p_{\rm miss}$ region of $1p_{3/2}$. This reduces the difference in the probed effective densities between the two shells (see Fig. 1).
@@ III.1 Polarization transfer
The calculations were performed on an event-by-event basis. Using each event’s kinematics parameters ensures a full match of the calculation to the experimental kinematics acceptance. It also allowed us to extract per-bin averages of the calculated observables. The original RDWIA program [9] was modified to include all 18 hadronic structure functions for the $\mathrm{A}(\vec{e},e^{\prime}\vec{p})$ reaction in the Born approximation [10]. The RDWIA calculations use the global democratic relativistic optical potential [11], relativistic bound-state wave functions obtained with the NL-SH parametrization [12], and free-proton EM FFs using the Bernauer parameterization [13].
@@ IV.1 Single-shell comparison with RDWIA
Figure 2: Single shell double ratios between the measured polarization transfer components and the ones calculated with RDWIA model from [9] as a function of the effective local nuclear density. From the fitted linear function and its $95\%$ confidence level band, a clear signature of a density-dependent effect with a negative slope can be seen.
@@ IV.2 Two-shell comparison with RDWIA
As inferred from the negative slope of the linear density dependence of the ratios shown in Fig. 2, the super ratios are on average below unity and tend to decrease with increasing density difference. We observe in Fig. 1 that between the $1s_{1/2}$ and $1p_{3/2}$ shells in ${}^{12}\mathrm{C}$ the density differs progressively with increase of $p_{\rm miss}$. The densities of the $2s_{1/2}$ and $1d_{3/2}$ shells in ${}^{40}\mathrm{Ca}$ stay much closer over the measured $p_{\rm miss}$ range. Accordingly, we observed deviations in the ratios between ${}^{12}\mathrm{C}$ shells being larger at high $p_{\rm miss}$ than at low $p_{\rm miss}$ and almost no effect present in the ratio between the $s$ and $d$ shells in ${}^{40}\mathrm{Ca}$ where densities are comparable. This suggests presence of density-dependent effect not yet included in RDWIA calculations.
@@ V In-medium proton modifications
The first investigation of polarization transfer sensitivity to possible density dependent in-medium modification of nucleon form factors was carried out by Kelly [1].This was followed by several polarization-transfer experiments. Measurements of the ${}^{4}\mathrm{He}(\vec{e}.e^{\prime}\vec{p}\,)$ reaction, performed both at MAMI [19, 20] and at JLab [21, 22], favored the calculations that included medium modifications models [23]. Two models were considered. The chiral quark soliton (CQS) model [24, 3] that mainly modifies valence quark contributions and the quark-meson coupling (QMC) model [4, 25]. However, the same data were later described by different calculations using free-proton EM FFs [26].
@@ VI Conclusions
Acknowledgements
We would like to thank the Mainz Microtron operators and the technical crew for the excellent operation of the accelerator. This work is supported by the Israel Science Foundation (Grant 951/19) of the Israel Academy of Arts and Sciences, by the Israel Ministry of Science, Technology and Spaces, by the PAZY Foundation (Grant 294/18), by the Deutsche Forschungsgemeinschaft (Collaborative Research Center 1044), by the U.S. National Science Foundation (PHY-2111050, PHY-2111442, and PHY-2239274), and by the United States-Israeli Binational Science Foundation (BS) as part of the joint program with the NSF (grant 2020742). We acknowledge the financial support from the Slovenian Research Agency (research core funding No. P1-0102).

### v3

@@ Abstract
The ratio of the transverse and longitudinal components of polarization transfer to protons in the quasi-elastic $(\vec{e},e^{\prime}\vec{p}\,)$ reaction, $P^{\prime}_{x}/P^{\prime}_{z}$, is sensitive to the proton’s electromagnetic form factor ratio, $G_{E}/G_{M}$. To explore density-dependent in-medium modifications, a comparison of polarization transfer ratios involving protons from distinct nuclear shells, each with different local nuclear densities, has been proposed. In this study, we present such comparisons between four shells, $1s_{1/2}$, $1p_{3/2}$ in ${}^{12}\mathrm{C}$ and $1d_{3/2}$, $2s_{1/2}$ in ${}^{40}\mathrm{Ca}$. In an effort to account for other many-body effects that may differ between shells, we use a state-of-the-art relativistic distorted-wave impulse-approximation (RDWIA) calculation and present the double ratios, $(P^{\prime}_{x}/P^{\prime}_{z})_{\rm Data}/(P^{\prime}_{x}/P^{\prime}_{z})_{\rm RDWIA}$ as well as the super ratios, $\left[(P^{\prime}_{x}/P^{\prime}_{z})_{\rm A}/(P^{\prime}_{x}/P^{\prime}_{z})_{\rm B}\right]_{\rm Data}/\left[(P^{\prime}_{x}/P^{\prime}_{z})_{\rm A}/(P^{\prime}_{x}/P^{\prime}_{z})_{\rm B}\right]_{\rm RDWIA}$, for chosen shells A and B, as a function of effective local nuclear densities. We find that double ratios for individual shells show a dependence on the probed effective nuclear densities. Studying the super ratios, we observed a systematic variation between pairs of higher- and lower-density shells.
@@ I Introduction
Polarization transfer to a proton bound in a nucleus has been suggested as a tool to observe in-medium modifications in the bound proton structure [1]. It is a part of a wider effort to understand the role of quarks and gluons in nuclei [2]. Some calculations introduce in-medium modifications and suggest nuclear-density-dependent changes of the bound nucleon electromagnetic (EM) form factors (FFs) [3, 4, 5]. We report on the first systematic search for nuclear density-dependent effects in the quasi-elastic $A(\vec{e},e^{\prime}\vec{p})$ reaction, which is sensitive to EM FFs.
@@
Density-dependent modifications are expected to be at the level of a few percent [5], which requires high statistical accuracy. ${}^{12}\mathrm{C}$ was suggested as a good nucleus for such studies since the effective local nuclear density experienced by the protons bound in the $1s_{1/2}$ shell is about twice the density for those in the $1p_{3/2}$ shell. Comparison of polarization ratios for protons from $1s_{1/2}$ and $1p_{1/2}$ shells in ${}^{12}\mathrm{C}$ has shown that the $p$ to $s$ double ratio was $1.15\pm 0.03$ [6]. Relativistic distorted-wave impulse-approximation (RDWIA) calculations only partially accounted for this deviation from unity, while the relativistic plane-wave impulse-approximation (RPWIA) did not predict this difference. However, once protons were compared at the same virtuality, the results became consistent with unity, $1.05\pm 0.05$, and so did RPWIA and RDWIA calculations [6]. We note that as seen in Fig. 2 of Ref. [6], in the virtuality overlap region, we are effectively comparing the low-$p_{\rm miss}$ region of $1s_{1/2}$ to the high-$p_{\rm miss}$ region of $1p_{3/2}$. This reduces the difference in the probed effective densities between the two shells (see Fig. 1).
@@ III.1 Polarization transfer
The calculations were performed on an event-by-event basis. Using each event’s kinematics parameters ensures a full match of the calculation to the experimental kinematics acceptance. It also allowed us to extract per-bin averages of the calculated observables. The original RDWIA program [9] was modified to include all 18 hadronic structure functions for the $\mathrm{A}(\vec{e},e^{\prime}\vec{p})$ reaction in the Born approximation [10]. The RDWIA calculations use the global democratic relativistic optical potential [11], relativistic bound-state wave functions obtained with the NL-SH parametrization [12], and free-proton EM FFs using the Bernauer parametrization [13].
@@ IV.1 Single-shell comparison with RDWIA
Figure 2: Single shell double ratios between the measured polarization transfer components and the ones calculated with the RDWIA model from [9] as a function of the effective local nuclear density. From the fitted linear function and its $95\%$ confidence level band, a clear signature of a density-dependent effect with a negative slope can be seen.
@@ IV.2 Two-shell comparison with RDWIA
As inferred from the negative slope of the linear density dependence of the ratios shown in Fig. 2, the super ratios are on average below unity and tend to decrease with increasing density difference. We observe in Fig. 1 that between the $1s_{1/2}$ and $1p_{3/2}$ shells in ${}^{12}\mathrm{C}$ the density differs progressively with increase of $p_{\rm miss}$. The densities of the $2s_{1/2}$ and $1d_{3/2}$ shells in ${}^{40}\mathrm{Ca}$ stay much closer over the measured $p_{\rm miss}$ range. Accordingly, we observed deviations in the ratios between ${}^{12}\mathrm{C}$ shells being larger at high $p_{\rm miss}$ than at low $p_{\rm miss}$ and almost no effect present in the ratio between the $s$ and $d$ shells in ${}^{40}\mathrm{Ca}$ where densities are comparable. This suggests the presence of a density-dependent effect not yet included in RDWIA calculations.
@@ V In-medium proton modifications
The first investigation of polarization transfer sensitivity to possible density dependent in-medium modification of nucleon form factors was carried out by Kelly [1].This was followed by several polarization-transfer experiments. Measurements of the ${}^{4}\mathrm{He}(\vec{e},e^{\prime}\vec{p}\,)$ reaction, performed both at MAMI [18] and at JLab [19, 20], favored the calculations that included medium modifications models [21]. Two models were considered. The chiral quark soliton (CQS) model [22, 3] that mainly modifies valence quark contributions and the quark-meson coupling (QMC) model [4, 23]. However, the same data were later described by different calculations using free-proton EM FFs [24].
@@ VI Conclusions
Acknowledgements.
We would like to thank the Mainz Microtron operators and the technical crew for the excellent operation of the accelerator. This work is supported by the Israel Science Foundation (Grant 951/19) of the Israel Academy of Arts and Sciences, by the Israel Ministry of Science, Technology and Spaces, by the PAZY Foundation (Grant No. 294/18), by the Deutsche Forschungsgemeinschaft (Collaborative Research Center 1044), by the U.S. National Science Foundation (Grants No. PHY-2111050, No. PHY-2111442, and No. PHY-2239274), and by the United States-Israeli Binational Science Foundation (BS) as part of the joint program with the NSF (Grant No. 2020742). We acknowledge the financial support from the Slovenian Research Agency (research core funding No. P1-0102).